BOBS POPCORN DELIGHT SHOP HAS TWO TYPES OF POPCORN. BUTTER POPCORN SELLS FOR 3.99 FOR A 12OZ BAG AND NON BUTTER POPCORN SELLS FOR 2.99 A 12 0Z BAG. HOW MANY BAGS OF EACH DOES BOB NEED TO MIX TO GET 50 12 OZ BAG TO PRODUCE A MIXTURE OF BOTH POPCORNS THAT SELLS FOR 3.49? ROUND EACH BAG TO THE NEAREST WHOLE NUMBER

Answer by checkley77(12844)

(3.99+2.99)/2=3.49
6.98/2=3.49
3.49=3.49
THEREFORE THE AMOUNT NEEDED (50 12 OZ BAGE) IS DIVIED EQUALLY BETWEEN THE 2 TYPES.
25 BAGS EACH.
PROOF
25*3.99+25*2.99=50*3.49
99.75+74.75=174.50
174.50=174.50